OpenStudy (anonymous):

Find all solutions to the equation csc x (2sinx-sqrt(2)) = 0 A. x= Pi/4 + kPi , where k is any positive integer. B. x= Pi/4 +2kPI and 7Pi/4 + 2kPI, where k is any positive integer C. x= 3pi/4 + 2Pik, where k is any positive integer. D. x = Pi/4 + 2kPi and 3Pi/4+ 2kPi, where k is any positive integer.

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

\[\csc x (2\sin x-\sqrt2) = 0\]

OpenStudy (anonymous):

thats the equation..

OpenStudy (anonymous):

\[{(2sinx-\sqrt2) \over \sin x} = 0\]\[{2\sin x \over \sin x}-{\sqrt2 \over \sin x}=0\]\[2-{\sqrt2 \over \sin x}=0\]

OpenStudy (anonymous):

\[2={\sqrt2 \over \sin x}\]

OpenStudy (anonymous):

\[2 \sin x=\sqrt2\]

OpenStudy (anonymous):

im not really seeing what you are saying..

OpenStudy (anonymous):

\[\sin x={\sqrt 2 \over 2}\]\[x={\pi \over 4},{3\pi \over 4}\]

OpenStudy (raden):

csc x (2sinx-sqrt(2)) = 0 for zeroes cscx=0 or 2sinx-sqrt(2) case I : cscx=0 1/sinx = 0 ----> no solution for x (it does not exis) case II : 2sinx-sqrt(2) = 0 2sinx = sqrt(2) sinx = 1/2 * sqrt(2) * for sinx = sin(pi/4) sinx=sin(2kpi+pi/4) x=2kpi+pi/4 ** for sinx=sin(3pi/4) sinx=sin(2kpi+3pi/4) x=2kpi+3pi/4

OpenStudy (anonymous):

so....

OpenStudy (raden):

D :p

OpenStudy (anonymous):

do you understand why it is D

OpenStudy (anonymous):

not really ..id like explaining...i don't really care for just answers..

OpenStudy (anonymous):

If you look at my work I prove that the two answers are pi/4 and 3pi/4. Before I explain the k stuff do you understand those steps?

OpenStudy (anonymous):

yeah i think so

OpenStudy (anonymous):

From the unit circle I knew those solutions. Also one full rotation around the circle is 2pi right?! so pi/4 + 2pi would be right back at the same spot, giving us the same answer. Then they added for any positive integer k because if this continued on for infinity, any positive integer would make just one more rotation 2pi, 4pi, 6pi, and so on. all giving the same value

OpenStudy (anonymous):

oh..okay

OpenStudy (anonymous):

Hope that helped